OpenStudy (barrelracer17):

For the Transformation T, write the T-1. T: (x, y) (x + 4, y + 3) T -1 (x, y) (x + 3, y + 2) (x - 4, y - 3) (x + ¼, y + )

OpenStudy (barrelracer17):

please help me ill reward you with a medal and a fan

OpenStudy (amoodarya):

t(x,y)=(X,Y)=(x + 4, y + 3) X=x+4 --->x=X-4 Y=y+3 --->y=Y-3 \[T^{-1}=(X-4,Y-3)\]